OpenStudy (aonz):

The total cost of producing x items per day is x^2/3 + 45x + 27 dollars, and the price per item at which each may be sold is 60− 12x dollars. Find an expression for the daily profit.

OpenStudy (aonz):

@oldrin.bataku help please

OpenStudy (anonymous):

So I can't help?

OpenStudy (anonymous):

lol @wio nice one.

OpenStudy (aonz):

umm anyone help please?

OpenStudy (anonymous):

You need to find the total revenue

OpenStudy (anonymous):

You have the revenue per item:\[ 60− 12x\]

OpenStudy (anonymous):

Multiply this by the number of items to get the revenue.

OpenStudy (anonymous):

@AonZ Does this help, or are you still lost?

OpenStudy (aonz):

It doesnt tell us the number of items sold?

OpenStudy (aonz):

btw this is part 1 of the question anyways

OpenStudy (anonymous):

What do you think \(x\) is?

OpenStudy (anonymous):

Firstly we should keep in mind that Profit is the difference between revenue and cost and shown like this:\[\bf P(x)=R(x)-C(x)\]Where 'P' is profit, 'R' is revenue, 'C' is cost. We are already given the cost function so we just need the revenue function.. We know that each item can be sold for \(\bf 60 -12x\) dollars hence the total revenue is given by the number of items multiplied by the selling price which gives us our revenue function as:\[\bf R(x)=x(60-12x)=60x-12x^2\]Now plug in C(x) to get P(x). Hence:\[\bf P(x)=R(x)-C(x)=60x-12x^2-\left( \frac{ 1 }{ 3 } x^2+45x+27)\right)\]\[\bf = 15x-\frac{ 37 }{ 3 }x^2-27\]

OpenStudy (anonymous):

@AonZ

OpenStudy (anonymous):

Well...

OpenStudy (aonz):

hmm answer in the back of my book doesnt match up with it :/

OpenStudy (aonz):

instead it has -5/6 x^2 + 15x - 27

OpenStudy (aonz):

nvm sorry you got it right

OpenStudy (anonymous):

maybe i made a computational mistake which i don't think i made but we can always double check

OpenStudy (anonymous):

=.=...

OpenStudy (aonz):

thanks to both of you

OpenStudy (anonymous):

@AonZ I hope you understood it